OpenStudy (piercetheveil47):

How could you eliminate the fractions to make solving 5/23m +12/23 =21/23 easier? A.) Divide both sides by 5/23 B.) Divide both sides of the equation by 23 C.) Multiply both sides of the equation by 23 D.) Subtract 12/23 from both sides of the equation

OpenStudy (piercetheveil47):

@Gin_Ichimaru hey! do you think you could help me with this??

OpenStudy (johnweldon1993):

If you multiply every term by 23 you eliminate the denominator... so multiplying both sides of the equation will make it \[\large \cancel{23} \times \frac{5}{\cancel{23}}m + \cancel{23} \times \frac{12}{\cancel{23}} = \cancel{23} \times \frac{21}{\cancel{23}}\] \[\large 5m + 12 = 21\]

OpenStudy (piercetheveil47):

thank you!!! @johnweldon1993